### v1

# Nodal Capacity Expansion Planning with Flexible Large-Scale Load Siting

Abstract
We propose explicitly incorporating large-scale load siting into a stochastic nodal power system capacity expansion planning model that concurrently co-optimizes generation, transmission and storage expansion. The potential operational flexibility of some of these large loads is also taken into account by considering them as consisting of a set of tranches with different reliability requirements, which are modeled as a constraint on expected served energy across operational scenarios. We implement our model as a two-stage stochastic mixed-integer optimization problem with cross-scenario expectation constraints. To overcome the challenge of scalability, we build upon existing work to implement this model on a high performance computing platform and exploit scenario parallelization using an augmented Progressive Hedging Algorithm. The algorithm is implemented using the bounding features of mpisppy, which have shown to provide satisfactory provable optimality gaps despite the absence of theoretical guarantees of convergence. We test our approach to assess the value of this proactive planning framework on total system cost and reliability metrics using realistic testcases geographically assigned to San Diego and South Carolina, with datacenter and direct air capture facilities as large loads.
[…]
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 and was funded by LDRD project 25-SI-007. We thank Gurobi for providing the academic license used to obtain the results provided in this work.
I Introduction
For over a decade, the electricity grid has been undergoing a transformation driven by the penetration of renewable, intermittent, and decentralized generation. While we continue to address the security and reliability difficulties posed by this transformation, the proliferation of large-scale electricity consumers, driven for example by datacenters as well as proposed direct air carbon capture (DAC) facilities, is further complicating an already taxing environment for grid planning and is a topic of increasing interest across academia, industry and policy makers [1, 2, 3].
In conventional planning methodologies, and consequently in existing Capacity Expansion Planning (CEP) software tools, it is customary to consider electricity load as a planning parameter, over which the planner has no control. The scale of the foreseen large-load increases challenges this view and pushes planners to take a more proactive role, seeking to influence load siting decisions as well as the possible co-location of on-site generation in addition to grid investments in generation, storage, and transmission. This approach takes advantage of the fact that subject to some constraints like access to low-latency areas for datacenters and geological suitability for carbon capture facilities, siting decisions can have a significant degree of flexibility. Such a framework requires updated software tools that can handle this new reality.
Besides siting flexibility, large loads can also provide value to the grid through operational flexibility. Demand Response (DR) programs exemplify this fact. CEP models that include expansion of DR as a decision variable have been proposed [4], but to the best of our knowledge, only in an aggregated zonal system, not as a nodal siting decision at the scale proposed here. The model we propose is a reliability-constrained CEP model, which is an active topic of research. In [5] a zonal CEP model is proposed where reliability is formulated as an expectation constraint across operational scenarios, which is then dualized to enable scenario decomposition, in the same vein of the method we propose here. This is only done for a single system-wide constraint instead of separately for each tranche of large load in the system like we do in this work. The idea of breaking load into tranches with differentiated reliability is well-known, and is implemented routinely at the consumer level through various mechanisms, including physically separated circuits (like in hospitals) and smart devices like controllable thermostats. Although we do not go into the details of ground-level implementation, the framework we propose here is inspired by the same concept.
It has been found that improving the spatial resolution of CEP models improves the quality of the investment plans obtained [6]. While computational challenges continue to exist, prior work has shown that High Performance Computing platforms can allow increasing both the number of operational scenarios and the spatial resolution of the power system considered while maintaining computational tractability [7].
In this paper, we address the challenges mentioned above by proposing a nodal power grid expansion planning model that explicitly considers sizing and siting large loads as part of the CEP co-optimization model, and considers their operational flexibility. The concrete contributions of our work are listed below.
Contributions
  • We introduce a two-stage mixed-integer stochastic optimization model for capacity expansion planning that explicitly considers large loads, like DAC and datacenters, part of the planning process, in addition to co-optimizing generation, transmission and storage expansion.
  • We propose a framework to represent the flexibility that large consumers can provide to the grid by breaking each large load into reliability tranches and adding expectation constraints to guarantee that the model captures both the limitations and the incentives of utilizing flexible large-load resources.
  • We develop an augmented Progressive Hedging Algorithm to decompose the expectation constraints added so that a scenario decomposition can be implemented in a parallel computing cluster, and implement it with mpisppy.
  • We illustrate through examples the value and promise of our model and solution technique.
Structure of this paper
Section II provides a description of the base CEP model and the modifications proposed to explicitly incorporate large load siting. Section III describes the solution methodology adopted to decompose the problem by scenario and corresponding solution in a parallel computing platform. Section IV presents and discusses some numerical tests performed; we close with some remarks in Section V.
II Model
II-A Base model
The basis for our CEP model is that of [8, 7]. It consists of a two-stage stochastic program that co-optimizes generation, transmission, and storage investments in the first stage and solves a multi-period optimal operation problem in the second stage. Second-stage suproblems, i.e., representative days, differ from each other in the hourly nodal demand and generation availability. This general framework is not uncommon in the literature of CEP [9, 10, 5].
Inverter-based generation and storage investments are modeled as continuous variables, turbine-based generation investments as integer variables (number of installed units), and transmission investments as binary variables. All investment costs are assumed to be linear in the size of the installed capacity. Storage levels are assumed to be cyclic. See [8] for a detailed discussion of the base model.
Next we provide a full, but compact description of the existing model with the minor changes made to incorporate load siting. In sections II-B and II-C we present more significant model extensions to handle the flexibility of these resources. A full list of the symbols used in this model is provided in Appendix A.
Objective
The objective is to minimize the total cost of expanding, maintaining, and operating the system, which we break into an investment cost $C^{\text{inv}}$ and an expected operation cost $\mathbb{E}_{\zeta}\left[C^{\text{op}}_{\omega}\right]$. The random variable $\zeta$ is used to represent in compact form all scenario-dependent data.
$C^{\text{inv}}=\sum_{\ell\in\mathcal{L}^{\star}}C^{\mathcal{L}f}_{\ell}x^{\mathcal{L}}_{\ell}+\\ +\sum_{b\in\mathcal{B}}\left(\sum_{g\in\mathcal{G}}C^{\mathcal{G}f}_{g}P^{\mathcal{G}}_{g}x^{\mathcal{G}}_{b,g}+\sum_{s\in\mathcal{S}}C^{\mathcal{S}f}_{s}x^{\mathcal{S}}_{b,s}+\sum_{d\in\mathcal{D}}C^{\mathcal{D}f}_{s}x^{\mathcal{D}}_{b,d}\right)$ (1a)
$\mathbb{E}_{\zeta}\left[C^{\text{op}}_{\omega}\right]=\sum_{\omega\in\Omega}\pi_{\omega}C^{\text{op}}_{\omega}=\sum_{\omega\in\Omega}\left(\pi_{\omega}\cdot 365\sum_{t\in\mathcal{T}}\sum_{b\in\mathcal{B}}\tau\cdot\Bigg[\right.\\ C^{\text{sh}}p^{\text{sh}}_{b,t,\omega}+\\ \left.\left.+\sum_{g\in\mathcal{G}}C^{\mathcal{G}v}_{g}p^{\mathcal{G}}_{b,g,t,\omega}+\sum_{s\in\mathcal{S}}C^{\mathcal{S}v}_{s}p^{\mathcal{S}\text{-dch}}_{b,s,t,\omega}+\sum_{d\in\mathcal{D}}C^{\mathcal{D}v}_{d}p^{\mathcal{D}}_{b,d,t,\omega}\right]\right)$ (1b)
The operation cost during each representative horizon $\omega$ is multiplied by 365 so that all costs are in \$/year.
Decision variables
Investment variables
The domains of the investment variables are those of (2). For large-scale loads, we only consider the case where investment variables must be integer multiples of a unit size. The model could easily be extended to also consider continuous load types.
$x^{\mathcal{G}}_{b,g}\in\mathbb{Z}^{+}~\forall b\in\mathcal{B},\forall g\in\mathcal{G}^{\mathbb{Z}}$ (2a)
$x^{\mathcal{G}}_{b,g}\in\mathbb{R}^{+}~\forall b\in\mathcal{B},\forall g\in\mathcal{G}^{\mathbb{R}}$ (2b)
$x^{\mathcal{S}}_{b,s}\in\mathbb{R}^{+}~\forall b\in\mathcal{B},\forall s\in\mathcal{S}$ (2c)
$x^{\mathcal{D}}_{b,d}\in\mathbb{Z}^{+}~\forall b\in\mathcal{B},\forall d\in\mathcal{D}$ (2d)
$x^{\mathcal{L}}_{\ell}\in\left\{0,1\right\}$$\forall\ell\in\mathcal{L}^{\star}$ (2e)
Operation variables
All operation variables are continuous (we do not include unit commitment or ramping in the second-stage problem). All operation variables except branch flows are non-negative.
Constraints
Construction constraints
We consider construction limits for new generation and storage resources. Limits are considered for each generation, storage, and load type at each bus (3). Limits across buses or types were omitted here but can easily be added to the model.
$P^{\mathcal{G}}_{g}x^{\mathcal{G}}_{b,g}+X^{\mathcal{G}}_{b,g}\leq K^{\mathcal{B}\mathcal{G}}_{b,g}~\forall b\in\mathcal{B},g\in\mathcal{G}$ (3a)
$x^{\mathcal{S}}_{b,s}+X^{\mathcal{S}}_{b,s}\leq K^{\mathcal{B}\mathcal{S}}_{b,s}~\forall b\in\mathcal{B},s\in\mathcal{S}$ (3b)
$x^{\mathcal{D}}_{b,d}\leq K^{\mathcal{B}\mathcal{D}}_{b,d}~\forall b\in\mathcal{B},d\in\mathcal{D}$ (3c)
Physical limits of generation and storage
The output of each generator cannot exceed the available power, which is dictated by the installed capacity and the availability of intermittent resources (4a). Similarly, the charge and discharge of storage facilities cannot exceed the installed power conversion capacity (4b),(4c), and the availability of storage must be respected (4d).
$~\forall b\in\mathcal{B},t\in\mathcal{T},\omega\in\Omega~:$
$p^{\mathcal{G}}_{b,g,t,\omega}\leq\alpha_{b,g,t,\omega}\left(X^{\mathcal{G}}_{b,g}+P^{\mathcal{G}}_{g}x^{\mathcal{G}}_{b,g}\right)~\forall g\in\mathcal{G}$ (4a)
$p^{\mathcal{S}\text{-ch}}_{b,s,t,\omega}\leq x^{\mathcal{S}}_{b,s}+X^{\mathcal{S}}_{b,s}~\forall s\in\mathcal{S}$ (4b)
$p^{\mathcal{S}\text{-dch}}_{b,s,t,\omega}\leq x^{\mathcal{S}}_{b,s}+X^{\mathcal{S}}_{b,s}~\forall s\in\mathcal{S}$ (4c)
$p^{\mathcal{S}}_{b,s,t,\omega}\leq U^{\mathcal{S}}_{s}\left(x^{\mathcal{S}}_{b,s}+X^{\mathcal{S}}_{b,s}\right)~\forall s\in\mathcal{S}$ (4d)
Energy storage
The change in energy storage level at each facility is driven by its charge and discharge (5a). To avoid end-of-horizon effects, we use the last period of the horizon as the initial storage state (5b). A constraint to impede simultaneous charging and discharging is omitted, as found to be unnecessary in previous work [10, 7].
$~\forall b\in\mathcal{B},s\in\mathcal{S},\omega\in\Omega,t\in\mathcal{T}\setminus\left\{\left|T\right|\right\}:$
$p^{\mathcal{S}}_{b,s,t,\omega}=p^{\mathcal{S}}_{b,s,t-1,\omega}+\tau\left(\eta^{\mathcal{S}\text{-ch}}_{s}p^{\mathcal{S}\text{-ch}}_{b,s,t,\omega}-p^{\mathcal{S}\text{-dch}}_{b,s,t,\omega}\right)$ (5a)
$p^{\mathcal{S}}_{b,s,0,\omega}=p^{\mathcal{S}}_{b,s,\left|\mathcal{T}\right|-1,\omega}+\tau\left(\eta^{\mathcal{S}\text{-ch}}_{s}p^{\mathcal{S}\text{-ch}}_{b,s,t,\omega}-p^{\mathcal{S}\text{-dch}}_{b,s,t,\omega}\right)$ (5b)
Load shedding
Shedded load cannot exceed demand (6). Note this excludes the large loads that are part of the planning process.
$p^{\text{sh}}_{b,t,\omega}\leq D_{b,t,\omega}~\forall b\in\mathcal{B},t\in\mathcal{T},\omega\in\Omega$ (6)
Power balance and power flow
We consider the standard $b\theta$ formulation of power flow (7a), with its corresponding big-M version for candidate transmission lines (7b). The energy balance is ensured at each individual bus (8).
$f_{\ell,t,\omega}=b_{\ell}(\theta_{o(\ell)}-\theta_{d(\ell)})~\forall\ell\in\mathcal{L}^{\dagger},t\in\mathcal{T},\omega\in\Omega$ (7a)
$M(1-x^{\mathcal{L}}_{\ell})\leq f_{\ell,t,\omega}-b_{\ell}(\theta_{o(\ell)}-\theta_{d(\ell)})\\ \leq M(1-x^{\mathcal{L}}_{\ell})~\forall\ell\in\mathcal{L}^{\star},t\in\mathcal{T},\omega\in\Omega$ (7b)
$\forall b\in\mathcal{B},t\in\mathcal{T},\omega\in\Omega:\\ \sum_{g\in\mathcal{G}}p^{\mathcal{G}}_{b,g,t,\omega}+\sum_{s\in\mathcal{S}}\left(\eta^{\mathcal{S}\text{-dch}}_{s}p^{\mathcal{S}\text{-dch}}_{b,s,t,\omega}-p^{\mathcal{S}\text{-ch}}_{b,s,t,\omega}\right)\\ ~~-\sum_{\ell\in\mathcal{L}:o(\ell)=b}f_{\ell,t,\omega}+\sum_{\ell\in\mathcal{L}^{\circ}:d(\ell)=b}f_{\ell,t,\omega}\\ ~~+p^{\text{sh}}_{b,t,\omega}-\sum_{d\in\mathcal{D}}p^{\mathcal{D}}_{b,d,t,\omega}=D_{b,t,\omega}$ (8)
Transmission limits
Limits for existing branches are enforced by (9a), while (9b) ensures them for built branches, as well as that unbuilt branches have no flow.
$-F_{\ell}\leq f_{\ell,t,\omega}\leq F_{\ell}~\forall\ell\in\mathcal{L}^{\dagger},t\in\mathcal{T},\omega\in\Omega$ (9a)
$-F_{\ell}x^{\mathcal{L}}_{\ell}\leq f_{\ell,t,\omega}\leq F_{\ell}x^{\mathcal{L}}_{\ell}~\forall\ell\in\mathcal{L}^{\star},t\in\mathcal{T},\omega\in\Omega$ (9b)
II-B Incorporating flexibility of large loads
Large loads that are included in the planning process are assumed to be composed of tiers with different reliability requirements. Each tier is defined by the fraction of total demand that it encompasses and its required reliability level, expressed as the expected capacity factor it must have. The scheme adopted to represent this is illustrated in Figure 1. Each tier $k=1,2,\dots,\left|\mathcal{K}\right|$ is defined by a pair $(u_{k},\phi_{k})$, with $u_{k},\phi_{k}\in\left[0,1\right]$. Tier $k$ consists of a fraction $u_{k}-u_{k-1}$ (with $u_{0}=0$) of the total installed capacity and must be served with a capacity factor of at least $\phi_{k}$.
Figure 1: Approach used to represent flexibility of large loads. Loads are assumed to be composed of tiers $k=1,\dots,\left|\mathcal{K}\right|$, with tier $k$ consisting of a fraction $u_{k}-u_{k-1}$ of total load, and requiring expected service of at least $\phi_{k}$.
Without loss of generality, we may sort the tiers by level of reliability and impose that $\phi_{k}$ be non-increasing. Note we must also have $u_{\left|\mathcal{K}\right|}=1$. Two special single-tier cases of worth are a completely inflexible load, given by $(u_{1},\phi_{1})=(1,1)$ and a fully flexible load, given by $(u_{1},\phi_{1})=(1,0)$. Finally, note that for simplicity we consider that all load types have the same number of tiers $\left|\mathcal{K}\right|$. This is without loss of generality since we can always add dummy tiers $(0,1)$ to any load type.
Flexibility tier constraints
Consistency of flexibility tiers is ensured by (10) and (11), while (12) ensures that the required reliability is respected.
$\forall b\in\mathcal{B},d\in\mathcal{D}:$
$p^{\mathcal{D}\mathcal{K}}_{b,d,k,t,\omega}\leq u_{d,k}x^{\mathcal{D}}_{b,d}~,\forall k\in\mathcal{K},t\in\mathcal{T},\omega\in\Omega$ (10)
$\sum_{k\in\mathcal{K}}p^{\mathcal{D}\mathcal{K}}_{b,d,k,t,\omega}=p^{\mathcal{D}}_{b,d,t,\omega}~,t\in\mathcal{T},\omega\in\Omega$ (11)
$\sum_{\omega\in\Omega}\pi_{\omega}\sum_{t\in\mathcal{T}}\tau\cdot p^{\mathcal{D}\mathcal{K}}_{b,d,k,t,\omega}\geq\phi_{d,k}u_{d,k}\left|\mathcal{T}\right|x^{\mathcal{D}}_{b,d}$ (12)
II-C Incentivizing construction and operation of large loads
Note that because the objective is cost minimization, it would be trivially optimal to not build any large loads and thus avoid the associated costs unless the model includes some incentive for their construction and operation. We discuss next different ways in which that can be addressed here.
II-C1 Mandated expansion
Perhaps the most direct and obvious way of incentivizing construction is by mandating it via a constraint like (13) below. This could be the case when system-wide expansion plans already exist, which could come from emissions policy in the case of DAC facilities, or expansion plans of private developers in the case of datacenters.
$\sum_{b\in\mathcal{B}}x^{\mathcal{D}}_{b,d}\geq M_{d}\forall d\in\mathcal{D}$ (13)
Note that if the type in question is fully inflexible, i.e. $(u_{d,1},\phi_{d,1})=(1,1)$, this mandate also provides a sufficient incentive for serving the load sites built.
II-C2 Incentive to produce output
For flexible types, an additional incentive must be given to utilize flexible resources; otherwise, it would be trivially optimal to only utilize each tier to its minimum service level required.
Via expectation constraints
We can provide such an incentive via expectation constraints imposing some bound on the output of those large loads, which could take the form of constraint (14).
$\sum_{\omega\in\Omega}\pi_{\omega}\left(q\cdot p^{\mathcal{G}}+r\cdot p^{\mathcal{D}}\right)\leq E$ (14)
The interpretation of vectors $q$ and $r$, and scalar $E$, differs depending on the type of load. For the case of carbon capture, $q$ could encode emission factors for each generation type, and $r$ could encode capture factors for each DAC type, so that $E$ would represent a net emissions target that may be set by policy. For the case of datacenters, $q=0$ and $r<0$ encodes how consumed energy translates to a metric of interest, with $E<0$ and E-E being the minimum threshold for that metric, for example server availability, or total training core-hours of AI models.
Note that if the threshold $E$ for the metric of interest has a non-trivial value (not 0), an expectation constraint in the style of (14) also provides an incentive for expansion, so that a mandate would be unnecessary.
Via negative costs
It is also possible to incentivize the utilization of flexible resources by encoding the economic value of their output into the objective, by substracting some reward term $r_{d}$ it from its variable cost, such that $r_{d}+C^{\mathcal{D}v}_{d}<0$. Presumably, the resulting value of $r_{d}-C^{\mathcal{D}v}_{d}$ would be in an intermediate range among the levelized cost of energy of the different generation types considered for expansion. Otherwise, the solution would trivially be to always or never serve that load type. In our experience, even in this case, having a net negative cost with a free sizing decision $x^{\mathcal{D}}_{d}$ leads to extreme solutions where load sites and generation capacity are built to their allowed maximum. This is a consequence of our simple linear model not taking into account the diminishing economic returns of the output $p^{\mathcal{D}}_{d}$.
While this could be addressed by considering a piecewise linear reward instead of a single-block $r_{d}$, we consider that case out of the scope of this paper. Instead, we limit models where costs are made negative to cases where we only have a load siting (not sizing) problem, i.e., when constraint (13) takes an equality form.
III Solution approach
For instances of moderate size, the mixed-integer linear program described above can be solved in Extensive Form $(EF)$. This is, however not a scalable approach as the size of the problem increases in number of scenarios and/or power system size.
[…]
III-A Relaxing expectation constraints
Large instances will necessitate decomposition, so the expectation constraints introduced need to be reformulated. As is standard practice, we will dualize these constraints. Let $\mathfrak{C}$ be the union set of all expectation constraints defined in (12) and (14), indexed by $c$. We write each constraint $c\in\mathfrak{C}$ in the form (15), then introduce a slack variable $\sigma_{c,\omega}$ given by (16) and associate a Lagrangian multiplier $\lambda_{c}\geq 0$ with it. We use here abstract symbols $x$ and $y_{\omega}$ to refer to the vectors of all first-stage and second-stage variables respectively, and $e_{c}$, $f_{c}$ and $h_{c}$ to refer to a scalar and vectors of coefficients of appropriate size.
$\sum_{\omega\in\Omega}\pi_{\omega}\left(f_{c}\cdot x+h_{c}\cdot y_{\omega}\right)\geq e_{c}$ (15)
$\sigma_{c,\omega}=e_{c}-f_{c}\cdot x+h_{c}\cdot y_{\omega}$ (16)
Doing this allows introducing the Lagrangian relaxation of $(EF)$, which we call $(LR)(\lambda)$.
[…]
Through standard duality arguments, we can conclude that $(LR)(\lambda)$ provides a lower bound of $(EF)$. Moreover, if strong duality holds, the optimal solution of $(EF)$ can be obtained by solving the dual problem $\max_{\lambda\geq 0}LR(\lambda)$. Note that in general, problem $(EF)$ might not be feasible. Relaxing and dualizing the constraints in the manner exposed here also ensures feasibility, which is convenient for computational implementations.
III-B Decomposition approach for large instances
By construction, the second-stage variables in $(LR)$ are no longer coupled across scenarios other than by the non-anticipativity of first-stage variables $x$. This methodology can thus be combined with existing decomposition techniques for two-stage stochastic optimization programs, like the Progressive Hedging Algorithm (PHA).
Recall that in the PHA, a copy of each first-stage variable is created for each scenario, and an equality constraint making all copies equal to each other is relaxed and dualized. In addition to the conventional Lagrangian term added to the objective, a so-called proximal term is added, resulting in problem $(PHA)_{\omega}$ for each scenario $\omega\in\Omega$. A more detailed description of the PHA is considered out of the scope of this paper. We refer the interested reader for example to [11, 12].
$(PHA)_{\omega}({\lambda}):$~ $\min\pi_{\omega}\cdot\left(C^{inv}+C^{op}_{\omega}+\sum_{c\in\mathfrak{C}}\lambda_{c}\sigma_{c,\omega}\right$.
$~~\left.+{w}\cdot x_{\omega}+\rho\left(x_{\omega}-\overline{x}\right)^{2}\right)$
[…]
We propose to solve our problem via the augmented PHA described below.
1: Initialize $w_{\omega}^{0}\leftarrow 0~\forall\omega,\lambda_{c}\leftarrow\lambda^{0}_{c}\forall c\in\mathfrak{C}$ , $k\leftarrow 0$, choose $\rho,\beta>0$
2:repeat
3:  for each scenario $\omega$ do
4:   Solve scenario subproblem $(PHA)_{\omega}$to get $x_{\omega}^{k+1}$:
5:     $x_{\omega}^{k+1}=\arg\min_{x_{\omega}\in\mathcal{X}_{\omega}}\left\{C^{inv}+C^{op}_{\omega}\right$.
6:            $\left.w_{\omega}^{k}x_{\omega}+\frac{\rho}{2}\|x_{\omega}-\bar{x}^{k}\|^{2}\right$.
7:            $\left.+\sum_{c\in\mathfrak{C}}\lambda^{k}_{c}\sigma_{c,\omega}\right\}$
8:  end for
9:  Update avg: $\bar{x}^{k+1}=\sum_{\omega}\pi_{\omega}x_{\omega}^{k+1}$
10:  for each scenario $\omega$ do
11:   Update multipliers: $w_{\omega}^{k+1}=w_{\omega}^{k}+\rho(x_{\omega}^{k+1}-\bar{x}^{k+1})$
12:  end for
13:  for each expectation constraint $c\in\mathfrak{C}$ do
14:   Update avg: $\bar{\sigma}^{k+1}_{c}=\sum_{\omega}\pi_{\omega}\sigma_{c,\omega}^{k+1}~\forall c\in\mathfrak{C}$
15:   Update multipliers: $\lambda_{c}^{k+1}=\lambda_{c}^{k}+\beta\bar{\sigma_{c}}$
16:  end for
17:  $k\leftarrow k+1$
18:until convergence criterion
About convergence guarantees
In the convex case, standard strong duality arguments can be used to guarantee convergence of the algorithm for an adequate choice of the step sizes $\beta$ and $\rho$ (see for example [5, 13]). Because of the presence of integers and binaries in the first-stage variables, our problem is not convex and thus neither convergence nor optimality at convergence are guaranteed. Consequently, this augmented PHA is, as the original PHA itself, only a heuristic technique on our non-convex case. However, it can be a satisfactory heuristic if methods are leveraged to obtain valid upper and lower bounds that provide quality guarantees of the solution obtained.
III-B1 Implementation in mpisppy
We implement our model in Pyomo, and use mpisppy to run the PHA utilizing parallel computing for the different scenario subproblems. The mpisppy package [14, 15] allows implementation in a parallel computing cluster with minimal effort, as it takes care of the necessary communication within the computing cluster during the PHA update steps. To implement the subgradient update feature that we describe here, we implement an mpisppy “Extension” to handle communication between the different compute nodes so that the update steps can be performed.
Obtaining upper and lower bounds on the solution
A key feature of mpisppy is its ability to provide quality guarantees during the execution of the algorithm. The package allows dedicating certain groups of computing nodes, referred to as spokes, to obtaining lower and upper bounds simultaneously as the PHA executes. Lower bounds may be obtained by solving relaxations of the problem, while upper bounds are obtained by testing feasible candidates. We refer the interested reader to our previous work [7] for more details about the implementation of a CEP model in a parallel computing cluster using mpisppy.
Obtaining lower bounds
Note that for the model proposed here, the values of $\lambda$ and/or $\bar{\sigma}$ need to be communicated to the bounder spokes. Custom versions of the mpisppy Lagrangian and reduced costs spokes [15] were developed for this purpose.
Obtaining upper bounds
In the conventional PHA, obtaining an upper bound consists in evaluating all second-stage subproblems for a candidate first-stage solution. In mpisppy, this is done by fixing the values of all non-anticipative variables to that of the evaluated candidate and solving the resulting second-stage subproblems (possibly in parallel). To facilitate evaluating upper bounds before the values of the dual multipliers have stabilized, we write another mpisppy “Extension” to fix the values of all first-stage variables but continue iterating on the values of $\lambda$ for a number of extra iterations. Note that in general, as discussed above, it is not guaranteed that for any given first-stage solution, all expectation constraints can be satisfied, i.e. we don’t necessarily have complete recourse. Once the values of first-stage variables is fixed, the resulting problem is an LP, and thus convexity guarantees stabilization of the values of $\lambda$. However, it is not guaranteed that this stabilization induces $\bar{\sigma}_{c}=0~\forall c\in\mathfrak{C}$.
IV Numerical results
We illustrate the value of the approach proposed here with a few numerical tests that we describe and discuss in this section.
IV-A Test cases
We implement our tests on two sandbox systems: the standard IEEE 24-bus test case, overlayed onto the San Diego area, and the ACTIVSg 500-bus South Carolina test case [16].
Both systems were extended with capacity expansion technoeconomic data and timeseries of load and generation availability downscaled from earth system models as in [8]. Some details of these test systems are reported in [17] and will be available in a separate publication in preparation at the time of writing. We consider two types of large-scale loads that are included in the expansion plan: “datacenter” and “DAC”. Table I summarizes the characteristics of the types considered. We conduct our tests on 12 operational scenarios with different demand and generation availability. These are sandbox testcases that are intended to represent a reasonable power system, but not necessarily a real one.
TABLE I: Characteristics of large-scale loads considered in tests.
Type Unit Size Utilization Incentive Flexibility
DAC 100 ktCO2/y $\bullet$ Net zero CO2 emissions $\bullet$ 2MWh/tCO2 captured Inflexible; Full-flex; Mid-flex: $u=(0.5,0.75,1)$, $\phi=(1,0.5,0)$
Data- center 800MW $\bullet$ Mandated datacenter target $\bullet$ Economic value: $C^{\mathcal{D}v}_{d}=-4\text{\textdollar/MWh}$ Mid-flex: $u=(0.5,0.9,1)$, $\phi=(1,0.85,0)$
IV-B Experiments
We first conduct a series of runs solving the problem in Extensive Form (EF) on the 24-bus instances, to better illustrate the value of explicitly modeling the flexibility of large loads as proposed in this paper. A handful of test cases are proposed to illustrate the features included in this model, in two groups. Test cases Ia, Ib, and Ic assume a net zero emissions target with different assumptions of flexibility for DAC facilities. Test cases IIa, IIb and IIc take Ia as base case and assume one 800MW datacenter must be built. Case IIc includes the datacenter siting problem with a set of 5 candidate sites with potential co-sited natural gas generation. Case IIa assumes that the new datacenter is added after the planning process is over, so it assumes the same buildout of case Ic and adds one datacenter at the bus selected in case IIc. Case IIb does the same but also adds one natural gas generator co-located with the new datacenter. A summary of the experiments presented here is shown in Table II.
To test the decomposition approach, we also implement instances Ia, Ib, Ic and IIc on both test case systems and compare the performance of the EF and decomposition approaches.
TABLE II: Summary of illustrative tests conducted
Test Load type Description
Ia DAC Inflexible
Ib DAC Mid-flexible
Ic DAC Fully flexible
IIa DAC+DatCent No extra planning, 1 datacenter
IIb DAC+DatCent Co-sited generation; 1 datacenter
IIc DAC+DatCent Proactive planning; 1 datacenter
IV-C Results
Fig. 2 shows the obtained buildout and costs for the experiments described above. Because of space limitations, only the results of the 24 bus test case are shown here. In all cases, both the EF and decomposition approaches return feasible solutions, i.e. the emissions target and reliability requirements are all satisfied. Note that despite the lack of a DAC mandate, the zero emissions expectation constraints indeed incentivizes construction of DAC facilities. An additional test case, not reported in these Figures, defined a 75% CO2 emissions reduction target. The model chose to decarbonize generation, meeting the target without resorting to DAC construction, which is in line with previous studies [18]. Note that in the test cases Ia-Ic proposed, the cases with increased flexibility are relaxations of the less flexible cases and should thus have lower costs. This is indeed what is found (Fig. 2, Right). The buildout provides some insight into how these savings occur: as large loads become more flexible, the system can afford to install less new storage and natural gas generators. In addition to requiring more generation, a new datacenter also necessitates a significant amount of new transmission with respect to the other cases.
Fig. 3 shows the results of the experiments with datacenter deployment with different levels of integration with the planning process. We observe two key insights in these results. First, failing to adequately expand the grid for the addition of such a significant load can lead to production costs orders of magnitude higher (Fig. 3, Left) and to not being able to provide the reliability these facilicies require (Fig. 3, Right). Co-siting of generation can alleviate the reliability issue, but can come at the expense of higher operational costs and compromise reaching the emission goals set.
Admittedly, these are sandbox systems with a limited number of scenarios, but we believe these results provide support for continuing to develop planning tools that incorporate large loads, and pay particular attention to modeling their flexibility.
Refer to captionRefer to caption
Figure 2: Left: Resource buildout for each test case. Case IIa has by construction the same buildout as Ic. Case IIb has the same buildout plus one natural gas unit co-sited with the datacenter; both are omitted from this figure. Transmission capacity is obtained by summing the capacities of all selected candidate lines or transformers. Right: Total cost and CO2 emissions for each test case.
Refer to captionRefer to caption
Figure 3: Left: Costs and emissions for cases IIa, IIb, IIc. Note that costs are in a log scale. Right: Achieved and required reliability for each tranche of each large load site built. Different shades of the same color show different tranches of the same site.
Performance of the decomposition approach
All tests were conducted on the dane cluster at Lawrence Livermore National Laboratory. We use Gurobi to solve each MILP subproblem. For 24-bus test cases, we used 3 compute nodes and a time limit of 30 minutes; for 500-bus test cases, 4 compute nodes and a time limit of 5 hours. Table III provides a summary of the computational results of the decomposition approach proposed on the different cases tested. Final costs are normalized with respect to the solution obtained with the EF formulation.
For the 24-bus test case, a very good candidate solution is obtained for instances Ib and Ic. The results are already satisfactory, but the lower bound could be tightened more with more sophisticated bounding techniques, for example implementing other spokes besides the reduced costs and Lagrangian ones. This could also help tightening the lower bound of the cases Ia and IIc, but in these two cases, the solution candidates obtained are suboptimal. Obtaining a better candidate would necessitate an improved update methodology for the dual multipliers, with perhaps dynamic values for $\beta$ and $\rho$. Improving this approach will be undertaken in future work but is outside the scope of the current paper.
For the 500-bus test cases, results are not yet satisfactory, but the computational bottleneck has been identified and a solution is under development Note however that the feasible solutions obtained with the decomposition approach are orders of magnitude better than the ones returned by the EF solver within the same time limit. It is true that the decomposition method had access to more computational resources (4x as many), so the comparison is not necessarily fair, but illustrates the potential of the methodology for instances where solving an EF version of the problem is not viable, which is precisely the intended use of this tool.
TABLE III: Performance of the decomposition approach
24 Buses 500 Buses
Test Final Gap Final Costa Final Gap Final Costa,b
Ia 26% 112% 75.2% 0.48%
Ib 4.3% 101% 76.4% 0.53%
Ic 4.0% 101% 75.0% 0.46%
IIc 73% 108% 80.1% 0.58%
  • a
    Normalized w.r.t. optimal cost returned by the EF.
  • b
    For all 500-bus instances, the EF returned very poor incumbents, with $99\%$ MIP gap.
V Closing remarks
We have presented here a model that explicitly considers the siting and sizing of large loads as part of the planning process. Our numerical results illustrate that proactively including these resources can save orders of magnitude in total system costs compared to building co-located generation while satisfying the reliability requirements of large loads. Moreover, explicitly considering the flexibility these resources are able to offer can provide a useful signal to planners for an optimal expansion plan that minimizes collective costs while making sure that the socially desired expansion occurs.
The 500-bus tests performed exhibit the computational challenges that arise with system size. Our tests of the proposed decomposition approach are promising and support continuing work to improve the method towards an implementation of a real-sized test case.
Appendix A List of symbols in optimization model
A-A Nomenclature
Sets
$\mathcal{B}$: Set of all buses (nodes) in the network. Indexed by $b$ (or $i,m,n$ where noted).
$\mathcal{G}$: Set of generation types, indexed by $g$, among which:
  $\mathcal{G}^{\mathbb{Z}}$: Types modeled with integer variables (turbine-based).
  $\mathcal{G}^{\mathbb{R}}$: Types modeled with continuous variables (inverter-based).
$\mathcal{S}$: Set of storage types, indexed by $s$.
$\mathcal{L}$: Set of transmission branches, indexed by $\ell$.
$\mathcal{L}^{\dagger}$: Set of existing transmission branches.
$\mathcal{L}^{\star}$: Set of candidate transmission branches.
$\mathcal{D}$: Set of large-scale load types, indexed by $d$
$\mathcal{K}$: Set of interruptibility tiers of large loads, indexed by $k$
$\mathcal{T}$: Set of periods in a representative horizon. Indexed by $t$.
$\Omega$: Set of scenarios in the uncertainty set of the stochastic optimization problem. Indexed by $\omega$.
$\mathcal{G}^{\mathbb{Z}}$ and $\mathcal{G}^{\mathbb{R}}$ constitute a partition of $\mathcal{G}$, i.e. $\mathcal{G}^{\mathbb{Z}}\cap\mathcal{G}^{\mathbb{R}}=\emptyset$ and $\mathcal{G}^{\mathbb{Z}}\cup\mathcal{G}^{\mathbb{R}}=\mathcal{G}$.
Index maps
$o(\ell)$: Origin bus (also called ‘from’ bus) of branch $\ell$
$d(\ell)$: Destination bus (also called ‘to’ bus) of branch $\ell$
Parameters
Existing resources
$X^{\mathcal{G}}_{b,g}$
: existing generation of type $g$ at bus $b$, in MW.
$X^{\mathcal{S}}_{b,s}$: existing storage power conversion capacity of type $s$ at bus $b$, in MW.
Fixed costs
$C^{(\cdot)f}_{(\cdot)}$
: Annualized fixed costs (capital + fixed O&M) for generation, storage or large load investment (in \$/MWy)
$C^{\mathcal{L}f}_{\ell}$: Annualized fixed costs for transmission candidate (in \$/y)
Variable costs
$C^{(\cdot)v}_{(\cdot)}$
: Variable costs (fuel, consumables, + variable O&M) for generation, storage and large load investment (in \$/MWh)
$C^{\text{sh}}$: Cost of load shedding, in \$/MWh.
Operational parameters
$P^{\mathcal{G}}_{g}$
: capacity per unit of generator type $g$, in MW.
$U^{\mathcal{S}}_{s}$: Duration of storage type $s$, in hours.
$D_{b,t,\omega}$: demand at bus $b$ during period $t$ of scenario $\omega$, excluding large loads that are part of the planning process.
$\alpha_{b,g,t,\omega}$: fraction of generation capacity of type $g$ that is available at bus $b$ during period $t$ of scenario $\omega$.
$\eta^{\mathcal{S}\text{-ch}}$: power conversion efficiency for $s$ when charging.
$\eta^{\mathcal{S}\text{-dch}}$: power conversion efficiency for $s$ when discharging.
$u_{d,k},\phi_{d,k}$: pair defining expected service level of tier $k$ of large load type $d$.
$b_{\ell}$: susceptance of branch $\ell$, in p.u.
$F_{\ell}$: transmission capacity of branch $\ell$ in MW.
Construction and planning parameters
$K^{\mathcal{B}(\cdot)}_{b,(\cdot)}$
: maximum buildable capacity of specific generation, storage or large load at bus $b$, in MW.
Other model parameters
$\tau$
: length of each period $t$ in a representative time horizon, in h.
$\pi_{\omega}$: probability assigned to scenario $\omega$ in the stochastic optimization problem.
Decision variables
Investment variables
$x^{(\cdot)}_{b,(\cdot)}$
: new storage, generation or large load of type $(\cdot)$ at bus $b$.
$x^{\mathcal{L}}_{\ell}$: binary variable indicating whether candidate $\ell$ is built.
Operation variables
$p^{\mathcal{G}}_{b,g,t,\omega}$
: output of generator of type $g$ at bus $b$ during period $t$ of scenario $\omega$, in MW.
$p^{\mathcal{S}}_{b,s,t,\omega}$: energy stored (level) in storage facility of type $s$ at bus $b$ during period $t$ of scenario $\omega$, in MWh.
$p^{\mathcal{S}\text{-ch}}_{b,s,t,\omega}$: power input (charging) of storage facility of type $s$ at bus $b$ during period $t$ of scenario $\omega$, in MW.
$p^{\mathcal{S}\text{-dch}}_{b,s,t,\omega}$: power output (discharging) of storage facility of type $s$ at bus $b$ during period $t$ of scenario $\omega$, in MW.
$p^{\mathcal{D}\mathcal{K}}_{b,d,t,k,\omega}$: power consumed by tier $k$ of large load of type $d$ at bus $b$ during period $t$ of scenario $\omega$, in MW.
$p^{\text{sh}}_{b,t,\omega}$: load shed at bus $b$ during period $t$ of scenario $\omega$, in MW (excluding large loads that are part of the planning process).
$f_{\ell,t,\omega}$: power flow through branch $\ell$ during period $t$ of scenario $\omega$, in MW.
Auxiliary variables and expressions
$C^{\text{inv}}$
: Annualized total investment cost in \$/y.
$C^{\text{op}}_{\omega}$: Operation cost of scenario $\omega$ in \$/y.
$p^{\mathcal{D}}_{b,d,t,\omega}$: total power consumed by large load of type $d$ at bus $b$ during period $t$ of scenario $\omega$, in MW.
$\sigma_{(\cdot)}$: slack variable associated with an expectation constraint.
Notation
Unless otherwise specified, the vector is noted by omitting the corresponding index, e.g. $f=\left[f_{\ell}\right]_{\ell\in\mathcal{L}}$. To allow for a more compact description, $(\cdot)$ was used above as a wildcard in superscripts and subscripts and may be substituted by $\mathcal{G},g$, $\mathcal{S},s$ or $\mathcal{D},d$ as appropriate.